Researchers are interested in the relationship between depression and bacterial vaginosis among expecting mothers enrolled in a case-control study. Suppose the following computerized printout reports a summary table for the relationship.

Table of dep by bv

dep

bv

_case

control

Total

_exposed

35

89

124

nonexposed

45

503

548

Total

80

592

672

A) Calculate the prevalence of the exposure in the investigation.

B) Calculate the prevalence of the disease in the investigation.

C) Calculate the prevalence odds ratio depicting the relationship between the outcome and the exposure.

Choose the most appropriate option

Option 1

a. 12.5%

b. 24.45%

c. 3.25

Option 2

a. 11.9%

b. 18.45%

c. 4.40

Option 3   

a. 22.6%

b. 14.45%

c. 5.45
Answer #1

A) prevalence of the exposure in the investigation =80/672 =0.119 ~ 11.90%

B) prevalence of the disease in the investigation =(124/672)=0.1845 ~ 18.45%

C) prevalence odds ratio depicting the relationship between the outcome and the exposure

=(35/89)/(45/503)=4.40

option 2